Assume that your father is now 55 years old and plans to retire after 5 years from now. 

He is expected to live for another 15 years after retirement. He wants a fixed retirement 

income of Rs. 1,00,000 per annum. His retirement income will begin the day he retires, 

5 years from today, and then he will get 14 additional payments annually. He expects to 

earn a return on his savings @ 10% p.a., annually compounding. How much (to the 

nearest of rupee) must your father save today to meet his retirement goal?



1
Expert's answer
2021-05-25T14:23:56-0400

P = Amount required annually = 100000

n = 15 years

r = return = 10%


Amount required at retirement:

"P+\frac{P(1-(1+r)^{-n-1})}{r}=100000+\frac{100000(1-(1+0.1)^{-15-1})}{0.1}=836668.75"


Calculation of Annual savings:

n = 5 years

r = annual return = 10%

Let p is Annual Savings required, then:

"p((1+r)^n-1)\/r=836668.75"

"p((1+0.1)^5-1)\/0.1=836668.75"

"0.61051p=83668.75"

"p=137047.30"